# Viability Constraints on Baryogenesis in $f(R,L_{m},T)$ Gravity

### v1

@@ III.1.2 Radiation-dominated era ($\omega=\frac{1}{3}$)
$\frac{\eta_{B}}{s}\;\Big{|}_{\omega=\frac{1}{3}}=-\frac{4.81959\times 10^{-56}(-\beta-2\gamma-50.24)^{3/4}\left(1.08334\times 10^{84}\sqrt{\alpha}\sqrt{-\beta-2\gamma-50.24}+1.32923\times 10^{36}\right)}{\alpha^{3/4}(\beta+2.\gamma+50.24)}.$ (29)
@@
Fig. 2 illustrates the variation of BnER $(\frac{\eta_{B}}{s})$ versus the model parameter $\alpha$, considering three representative choices of $\beta=-1\times 10^{2}$, $\beta=-5\times 10^{2}$ and $\beta=-1\times 10^{3}$. Throughout this analysis, the parameter $\gamma$ is kept fixed at $\gamma=1$. The dashed red line in the figure indicates the observationally inferred value of baryon asymmetry, $\frac{\eta_{B}}{s}$. The plot depicted in Fig. 2, suggests that the theoretically calculated asymmetry ratio for the discussed model is consistent with observational constraints when $\alpha$ lies approximately within the range: ($1.5\times 10^{157},9.0\times 10^{157}$).
III.1.3 Stiff fluid-dominated era ($\omega=1$)
For stiff regime, BnER in Eq. (27) reduces to the following simplified expression:
@@ III.1.3 Stiff fluid-dominated era ($\omega=1$)
$\frac{\eta_{B}}{s}\;\Big{|}_{\omega=1}=\frac{2.11957\times 10^{-6}(-\beta-2\gamma-50.24)^{3/4}}{\alpha^{3/4}}-\frac{7.07692\times 10^{28}\sqrt[4]{-\beta-2\gamma-50.24}}{\alpha^{1/4}}.$ (30)
@@
Fig. 3 illustrates the variation of the BnER $(\frac{\eta_{B}}{s})$ against the parameter $\alpha$ for distinct $\beta$ parameter values of $\beta=-1\times 10^{2}$, $\beta=-5\times 10^{2}$ and $\beta=-1\times 10^{3}$. In this case, the parameter $\gamma$ is fixed at $\gamma=1$. A comparison between the theoretical curves and baryon asymmetry inferred from cosmological data, represented by the horizontal red dashed line in Fig. 3 suggests approximate value of $\alpha$ to lie in the interval ($9\times 10^{157},3\times 10^{158}$) and the result is in close agreement with observations.
III.1.4 Quintessence era ($1<\omega<-\frac{1}{3}$)
Fig. 4 depicts the dependence of the BnER $(\frac{\eta_{B}}{s})$ on $\omega$ for three contrasting $\alpha$ magnitudes, namely, $\alpha=3.56\times 10^{157}$, $\alpha=1.98\times 10^{157}$ and $\alpha=5.70\times 10^{156}$. The other model parameters $\beta$ and $\gamma$ are fixed at: $\beta=-1\times 10^{2}$ and $\gamma=1$ for analysis done in the case of quintessence era. The horizontal line (red-dashed) corresponds to the baryon asymmetry inferred from observational data. From Fig. 4, it is evident that the calculated imbalance ratio for the model is consistent with the observed baryon asymmetry in quintessence case.
@@ IV Generalized Gravitational Baryogenesis in $f(R,L_{m},T)$
Here, we extend into the generalized case by considering a more comprehensive $\mathcal{CP}$-violating interaction for the chosen $f(R,L_{m},T)$ model. The interaction term in gravitational baryogenesis is directly linked to ($R+L_{m}+T$), the generalized gravitational baryogenesis formulation introduces a coupling term proportional to the function itself, thereby enriching the dynamics of $\mathcal{CP}$-violation. The corresponding coupling term in this context in this context can thus be expressed as [24]:
@@
$\begin{split}\frac{\eta_{B}}{s}\;\Big{|}_{GGB}=-(4.48514\times 10^{-56}((\omega+1)(-\beta-2\gamma-50.24))^{3/4}(\sqrt{\alpha}(\omega(\beta+2\gamma+50.24)\times 10^{50}+(-\beta-2\gamma-50.24)\times 10^{50})\\ \hskip 0.0pt+(\gamma(2.8146\times 10^{84}\omega-9.382\times 10^{83})-9.382\times 10^{83}\beta)\sqrt{(\omega+1)(-\beta-2\gamma-50.24)}))/(\sqrt[4]{\alpha}\sqrt{\omega+1}(\beta+2.\gamma+50.24)).\end{split}$ (33)
To analyse the generalized gravitational baryogenesis for $f(R,L_{m},T)$ model, the parameters having fixed magnitudes in Eq. (33) are: $g_{*}=106$, $g_{b}\simeq 1$, $M_{*}=2\times 10^{18}GeV$, and the decoupling temperature is chosen as $T_{D}=2\times 10^{16}GeV$.
Figure 5: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ (generalized case) for the $f(R,L_{m},T)$ model under consideration in dust regime $(\omega=0)$ for three different pairs $\beta$ and $\gamma$ choices.
Figure 6: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ (generalized case) for the $f(R,L_{m},T)$ model under consideration in radiation era $(\omega=\frac{1}{3})$ for three different $\beta$ choices.
Figure 7: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ (generalized case) for the $f(R,L_{m},T)$ model under consideration in stiff regime $(\omega=1)$ for three different pairs $\beta$ and $\gamma$ choices.
Figure 8: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ (generalized case) for the $f(R,L_{m},T)$ model under consideration in quintessence era (three different $\omega$ values are plotted, Blue($\omega=-0.50$), Yellow($\omega=-0.65$) ,Green($\omega=-0.85$)) for three different pairs of $\beta$ and $\gamma$ choices.
IV.0.1 Matter-dominated era ($\omega=0$): Dust case
For matter-dominated scenario with $\omega=0$, the BnER in Eq. (33) reduces to:
$\frac{\eta_{B}}{s}\;\Big{|}_{GGB,\omega=0}=\frac{\sqrt[4]{-\beta-2\gamma-50.24}\left(8.91168\times 10^{-6}\sqrt{\alpha}\sqrt{-\beta-2\gamma-50.24}-4.20796\times 10^{28}\beta-4.20796\times 10^{28}\gamma\right)}{\sqrt[4]{\alpha}}.$ (34)
Fig. 5 presents the variation of BnER $(\frac{\eta_{B}}{s})_{ggb}$ on $\alpha$ in the context of generalized gravitational baryogenesis. In matter-dominated case, $\omega=0$, the ratio exhibits a pronounced sensitivity to the chosen combinations of $\beta$ and $\gamma$. The parameter pairs considered for the plot are: $(1.0\times 10^{2},-1.0\times 10^{2})$, $(1.2\times 10^{2},-1.2\times 10^{2})$ and $(1.4\times 10^{2},-1.4\times 10^{2})$. The horizontal dashed line in the plot denotes the observationally established value of baryon asymmetry, $\frac{\eta_{B}}{s}=9.42\times 10^{-11}$. From the Fig. 5, it is evident that the theoretical predictions reproduce the observed asymmetry when $\alpha$ lies within the approximate interval: ($2\times 10^{-26},1.0\times 10^{-25}$), suggesting close alignment with observational constrain.
IV.0.2 Radiation-dominated era ($\omega=\frac{1}{3}$)
@@
$\begin{split}\frac{\eta_{B}}{s}\;\Big{|}_{GGB,\omega=\frac{1}{3}}&=-(1.60653\times 10^{-56}(-\beta-2\gamma-50.24)^{3/4}((-3.25002\times 10^{84}\beta-3.73568\times 10^{68}\gamma)\sqrt{-\beta-2\gamma-50.24}\\ &\quad-3.98768\times 10^{36}\sqrt{\alpha}))/(\sqrt[4]{\alpha}(\beta+2.\gamma+50.24)).\end{split}$ (35)
@@
Fig. 6 displays the behaviour of the generalized BnER $(\frac{\eta_{B}}{s})_{ggb}$ in terms of $\alpha$ for distinct choices of $\beta$, namely $\beta=-6\times 10^{2}$, $\beta=-8\times 10^{2}$ and $\beta=-1\times 10^{3}$, while keeping $\gamma$ fixed at unity. The horizontal red dashed line in the figure corresponds to the baryon asymmetry derived from astronomical data. theoretical estimates in comparison with observational constrain concludes that the viable range of $\alpha$ lies approximately within the interval ($0.5\times 10^{169},9\times 10^{169}$), indicating strong consistency between the model outcomes and empirical data.
IV.0.3 Stiff fluid-dominated era ($\omega=1$)
When the cosmic background corresponds to the stiff case $\omega=1$, the expression for the BnER obtained from Eq. (33) simplifies as follows:
$\begin{split}\frac{\eta_{B}}{s}\;\Big{|}_{GGB,\omega=1}&=-(5.33376\times 10^{-56}(-\beta-2\gamma-50.24)^{3/4}(\sqrt{\alpha}(\beta+2\gamma+50.24)\times 10^{52}\\ &\quad+(2.65363\times 10^{84}\gamma-1.32682\times 10^{84}\beta)\sqrt{-\beta-2\gamma-50.24}))/(\sqrt[4]{\alpha}(\beta+2.\gamma+50.24)).\end{split}$ (36)
@@ IV.0.3 Stiff fluid-dominated era ($\omega=1$)
Fig. 7 presents the evolution of the asymmetry ratio $(\frac{\eta_{B}}{s})_{ggb}$ with respect to the free parameter $\alpha$. The outcome is highly sensitive to the parameters $\beta$ and $\gamma$ as a correlated binary set. In this study, three specific pairings were examined: $(-2.2\times 10^{2},-1.1\times 10^{2})$, $(-2.6\times 10^{2},-1.3\times 10^{2})$ and $(-3.0\times 10^{2},-1.5\times 10^{2})$. The horizontal dashed red line in the figure marks the baryon asymmetry deduced from cosmological observations, serving as a reference point. By comparing the plotted trends with this observational indicator, one finds that the parameter $\alpha$ falls consistently within the interval ($2.2\times 10^{-30},6.0\times 10^{-30}$), showing strong compatibility between theoretical estimates and empirical evidence.
IV.0.4 Quintessence era ($1<\omega<-\frac{1}{3}$)
Fig. 8 presents the variation of BnER $(\frac{\eta_{B}}{s})_{ggb}$ against $\alpha$. Due to the mathematical intricacy of the imbalance ratio expression for the generalized case in quintessence era, the asymmetry ratio cannot be directly expressed as a function of $\omega$. Therefore, the analysis is restricted to three distinct values of $\omega$ such that: $0.3>\omega>-1$. In quintessence era, $(\frac{\eta_{B}}{s})_{ggb}$ exhibit delicate dependency on the combined doublet $(\beta,\gamma)$. The choice of three such $(\beta,\gamma)$-pairs taken for plotting are: $(-2.0\times 10^{2},52)$, $(-2.5\times 10^{2},65)$ and $(-3.0\times 10^{2},78)$. As illustrated in Fig. 8, the theoretical estimates remain in agreement with the observed asymmetry values (marked by the red-dashed horizontal line) throughout the quintessence regime.

### v2

@@ III.1.1 Radiation-dominated era ($\omega=\frac{1}{3}$)
$\frac{\eta_{B}}{s}\;\Big|_{\omega=\frac{1}{3}}=-\frac{9.64\times 10^{-49}(-\beta-2\gamma-50.24)^{3/4}\left(3.39\times 10^{47}\times\sqrt{\alpha(-\beta-2\gamma-50.24)}+1.41\times 10^{14}\right)}{\alpha^{3/4}(\beta+2\gamma+50.24)}.$ (28)
@@
Fig. 1 illustrates the variation of BnER $(\frac{\eta_{B}}{s})$ versus the model parameter $\alpha$, considering three representative choices of $\beta=-1\times 10^{2}$, $\beta=-5\times 10^{2}$ and $\beta=-1\times 10^{3}$ for Model I. Throughout this analysis, the parameter $\gamma$ is kept fixed at $\gamma=1$. The dashed red line in the figure indicates the observationally inferred value of baryon asymmetry, $\frac{\eta_{B}}{s}$. The plot depicted in Fig. 1, suggests that the theoretically calculated asymmetry ratio for the discussed model is consistent with observational constraints.
III.2 Model II
For this case, we employ the $f(R,L_{m},T)$ model formulated as follows [30, 32]:
$f(R,L_{m},T)=\alpha R+\beta L_{m}^{2}+\gamma T^{2},$ (29)
here, $\alpha$, $\beta$ and $\gamma$ are free parameters of the model.
The energy density for Model II can be deduced in a similar manner as follows:
@@ III.2 Model II
$\rho=-\frac{2\alpha\xi^{2}}{\sqrt{2}\sqrt{t^{4}(\omega+1)\left(\alpha\xi^{2}(\beta-6\gamma\omega+2\gamma)+8\pi^{2}t^{4}(\omega+1)\right)}+4\pi t^{4}(\omega+1)}.$ (30)
The expression for decoupling time $t_{D}$ is obtained as:
$t_{D}=\frac{\sqrt{15}\left(\frac{2}{\pi}\right)^{3/4}\sqrt[4]{\alpha}\sqrt{\xi}}{\sqrt[4]{g_{*}T_{D}^{4}(\omega+1)\left(\pi g_{*}T_{D}^{4}(\beta-6\gamma\omega+2\gamma)-240\right)}}.$ (31)
Hence, the final expression of BnER for Model II is obtained as:
$\frac{\eta_{B}}{s}=\frac{15g_{b}(J(2-3\omega)-K)}{4\pi^{2}g_{*}M_{*}^{2}T_{D}},$ (32)
where,
$J$ $=\frac{8\alpha\left(C+\frac{\pi^{3/2}(D+G)}{40\sqrt{F}}\right)}{9(\omega+1)^{2}\left(\frac{3200\alpha}{\pi^{2}Bg_{*}T_{D}^{4}(\omega+1)^{2}}+\frac{40\sqrt{F}}{\pi^{3/2}}\right)^{2}},\hskip 142.26378ptK=\frac{\pi^{9/4}\left(Bg_{*}T_{D}^{4}(\omega+1)\right)^{3/4}\left(1-\frac{4}{3(\omega+1)}\right)}{5\ 2^{3/4}\sqrt{5}\alpha^{3/4}\sqrt{\frac{1}{\omega+1}}},$
$G$ $=\frac{160\ 2^{3/4}\sqrt{5}\alpha^{3/4}\sqrt{\frac{1}{\omega+1}}\left(\frac{4\alpha A}{9(\omega+1)^{2}}+\frac{6400\alpha}{\pi Bg_{*}T_{D}^{4}(\omega+1)^{2}}\right)}{\pi^{9/4}\left(Bg_{*}T_{D}^{4}(\omega+1)\right)^{3/4}},\hskip 79.6678ptF=\frac{\alpha\left(\frac{4\alpha A}{9(\omega+1)^{2}}+\frac{6400\alpha}{\pi Bg_{*}T_{D}^{4}(\omega+1)^{2}}\right)}{Bg_{*}T_{D}^{4}(\omega+1)^{2}},$
$D$ $=\frac{1024000\ 2^{3/4}\sqrt{5}\alpha^{7/4}\left(\frac{1}{\omega+1}\right)^{3/2}}{\pi^{13/4}\left(Bg_{*}T_{D}^{4}(\omega+1)\right)^{7/4}},\hskip 165.02606ptC=\frac{640\ 2^{3/4}\sqrt{5}\alpha^{3/4}\sqrt{\frac{1}{\omega+1}}}{\pi^{5/4}\left(Bg_{*}T_{D}^{4}(\omega+1)\right)^{3/4}},$
$B$ $=\left(\pi Ag_{*}T_{D}^{4}-240\right),\hskip 227.62204ptA=\left(\beta-6\gamma\omega+2\gamma\right).$
For Eq. (32), the set of constant parameters is taken to be identical to that used in Model I.
Figure 2: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ for the $f(R,L_{m},T)$ model II under consideration in radiation era $(\omega=\frac{1}{3})$ for three different $\beta$ choices.
III.2.1 Radiation-dominated era ($\omega=\frac{1}{3}$)
For radiation-dominated regime, BnER in Eq. (32) reduces to the following simplified expression:
$\frac{\eta_{B}}{s}\;\Big|_{\omega=\frac{1}{3}}=\frac{\pi^{3/4}\alpha^{3/4}g_{b}\sqrt[4]{g_{*}T_{D}^{4}\left(\pi\beta g_{*}T_{D}^{4}-240\right)}}{2\sqrt{5}6^{3/4}g_{*}M_{*}^{2}T_{D}\sqrt{\frac{\alpha^{2}\left(\pi\beta g_{*}T_{D}^{4}-120\right)^{2}}{g_{*}^{2}T_{D}^{8}\left(\pi\beta g_{*}T_{D}^{4}-240\right)^{2}}}}.$ (33)
@@ III.2.1 Radiation-dominated era ($\omega=\frac{1}{3}$)
Fig. 2 illustrates the variation of the BnER $(\frac{\eta_{B}}{s})$ against the parameter $\alpha$ for distinct $\beta$ parameter values of $\beta=1\times 10^{-28}$, $\beta=5\times 10^{-28}$ and $\beta=1\times 10^{-27}$ for Model II. In this case, the parameter $\gamma$ is fixed at $\gamma=1$. A comparison between the theoretical curves and BnER inferred from cosmological data, represented by the horizontal red dashed line in fig. 2, suggests close agreement with observations.
@@ IV Generalized Gravitational Baryogenesis in $f(R,L_{m},T)$
Here, we extend into the generalized case by considering a more comprehensive $\mathcal{CP}$-violating interaction for the chosen gravity models. The interaction term in gravitational baryogenesis is directly linked to ($R+L_{m}+T$), while, the generalized gravitational baryogenesis formulation introduces a coupling term proportional to the function itself, thereby enriching the dynamics of $\mathcal{CP}$-violation. The corresponding coupling term in this context in this context can thus be expressed as [26]:
@@ IV.1 Model I
$\begin{split}\frac{\eta_{B}}{s}\;\Big|_{GGB}&=\left(45\times\sqrt{2\alpha}(2\xi-1)\sqrt{g_{*}(\omega+1)(-\beta-2\gamma-16\pi)T_{D}^{4}}-2\sqrt{15}\pi g_{*}T_{D}^{4}(\beta-3\gamma\omega+\gamma)\right)\\ &\quad\times\frac{g_{b}\sqrt[4]{g_{*}(\omega+1)(-\beta-2\gamma-16\pi)T_{D}^{4}}}{4\times 30^{3/4}\sqrt{\pi}\sqrt[4]{\alpha}g_{*}M_{*}^{2}\sqrt{\xi}T_{D}}.\end{split}$ (36)
To analyse the generalized case of gravitational baryogenesis for $f(R,L_{m},T)$ model, the parameters having fixed magnitudes in Eq. (36) are: $g_{*}=106$, $g_{b}\simeq 1$, $M_{*}=2\times 10^{18}GeV$, and the decoupling temperature is chosen as $T_{D}=2\times 10^{9}GeV$.
Figure 3: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ (generalized case) for the $f(R,L_{m},T)$ model I under consideration in radiation era $(\omega=\frac{1}{3})$ for three different $\beta$ choices.
IV.1.1 Radiation-dominated era ($\omega=\frac{1}{3}$)
@@
$\begin{split}\frac{\eta_{B}}{s}\;\Big|_{GGB,\omega=\frac{1}{3}}&=\frac{(2.27\times 10^{-15}+9.03\times 10^{-17}\gamma)\gamma}{\sqrt[4]{\alpha}(-\beta-2\gamma-50.24)^{3/4}}-\frac{0.33\beta\times(-\beta-2\gamma-50.24)^{1/4}}{\sqrt[4]{\alpha}}-\frac{1.36\times 10^{-34}\times\sqrt[4]{\alpha}}{(-\beta-2\gamma-50.24)^{1/4}}.\end{split}$ (37)
@@
Fig. 3 displays the behaviour of the generalized BnER $(\frac{\eta_{B}}{s})_{ggb}$ in terms of $\alpha$ for distinct choices of $\beta$, namely $\beta=-6\times 10^{2}$, $\beta=-8\times 10^{2}$ and $\beta=-1\times 10^{3}$, while keeping $\gamma$ fixed at unity. The horizontal red dashed line in the figure corresponds to the baryon asymmetry derived from astronomical data. Theoretical estimates, in comparison with observational constraints, indicate a strong consistency between the model outcomes and empirical data.
IV.2 Model II
For Model II, the expression of BnER is obtained in the following form:
$\frac{\eta_{B}}{s}\;\Big|_{GGB}=\frac{15g_{b}\left(\beta J+\gamma(J)(3\omega-1)^{2}-K\right)}{4\pi^{2}g_{*}M_{*}^{2}T_{D}},$ (38)
where,
$J$ $=\frac{128\alpha^{2}\left(C+\frac{\pi^{3/2}(D+G)}{40\sqrt{F}}\right)}{81(\omega+1)^{4}\left(\frac{3200\alpha}{\pi^{2}Bg_{*}T_{D}^{4}(\omega+1)^{2}}+\frac{40\sqrt{F}}{\pi^{3/2}}\right)^{3}},\hskip 102.43008ptK=\frac{\pi^{9/4}\sqrt[4]{\alpha}\left(1-\frac{4}{3(\omega+1)}\right)\left(Bg_{*}T_{D}^{4}(\omega+1)\right)^{3/4}}{5\ 2^{3/4}\sqrt{5}\sqrt{\frac{1}{\omega+1}}},$
$G$ $=\frac{160\ 2^{3/4}\sqrt{5}\alpha^{3/4}\sqrt{\frac{1}{\omega+1}}\left(\frac{4\alpha A}{9(\omega+1)^{2}}+\frac{6400\alpha}{\pi Bg_{*}T_{D}^{4}(\omega+1)^{2}}\right)}{\pi^{9/4}\left(Bg_{*}T_{D}^{4}(\omega+1)\right)^{3/4}},\hskip 51.21504ptF=\frac{\alpha\left(\frac{4\alpha A}{9(\omega+1)^{2}}+\frac{6400\alpha}{\pi Bg_{*}T_{D}^{4}(\omega+1)^{2}}\right)}{Bg_{*}T_{D}^{4}(\omega+1)^{2}},$
$D$ $=\frac{1024000\ 2^{3/4}\sqrt{5}\alpha^{7/4}\left(\frac{1}{\omega+1}\right)^{3/2}}{\pi^{13/4}\left(Bg_{*}T_{D}^{4}(\omega+1)\right)^{7/4}},\hskip 136.5733ptC=\frac{640\ 2^{3/4}\sqrt{5}\alpha^{3/4}\sqrt{\frac{1}{\omega+1}}}{\pi^{5/4}\left(Bg_{*}T_{D}^{4}(\omega+1)\right)^{3/4}},$
$B$ $=\left(\pi Ag_{*}T_{D}^{4}-240\right),\hskip 199.16928ptA=\left(\beta-6\gamma\omega+2\gamma\right).$
The fixed parameters utilized in Eq. (38) follow the identical numerical values as those employed in earlier cases.
Figure 4: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ (generalized case) for the $f(R,L_{m},T)$ model II under consideration in radiation era $(\omega=\frac{1}{3})$ for three different $\beta$ choices.
IV.2.1 Radiation-dominated era ($\omega=\frac{1}{3}$)
For radiation dominant regime, the expression for the BnER obtained from Eq. (38) simplifies as follows
$\frac{\eta_{B}}{s}\;\Big|_{GGB,\omega=\frac{1}{3}}=\frac{\pi^{11/4}\alpha^{11/4}\beta g_{b}P^{5/4}\left(28800\alpha+\pi\alpha\beta P+240P\sqrt{Q}\right)}{30\sqrt{5}6^{3/4}g_{*}M_{*}^{2}\sqrt{Q}T_{D}\left(120\alpha+P\sqrt{Q}\right)^{3}},$ (39)
where,
$P$ $=\Big(g_{*}T_{D}^{4}\left(\pi\beta g_{*}T_{D}^{4}-240\right)\Big),\hskip 113.81102ptQ=\left(\frac{\alpha^{2}\left(\pi\beta g_{*}T_{D}^{4}-120\right)^{2}}{g_{*}^{2}T_{D}^{8}\left(\pi\beta g_{*}T_{D}^{4}-240\right)^{2}}\right).$
@@ IV.2.1 Radiation-dominated era ($\omega=\frac{1}{3}$)
Fig. 4 shows how the imbalance ratio $(\frac{\eta_{B}}{s})_{ggb}$ evolves with the free parameter $\alpha$. The outcome is sensitive to the values of the $\beta$ parameter while $\gamma$ is fixed at unity. In fig. 4, three specific $\beta$ values are taken, $\beta=4\times 10^{-36}$, $\beta=6\times 10^{-36}$ and $\beta=8\times 10^{-36}$. The horizontal dashed red line in the figure marks the baryon asymmetry deduced from cosmological observations, serving as a reference point. By comparing the plotted graphs with this observational indicator, one finds strong compatibility between theoretical estimates and empirical evidence.